OpenStudy (anonymous):

prove: sinh^2(x)=-1+cosh(2x)/2

zepdrix (zepdrix):

So convert to exponentials and work it out! :)\[\Large\rm \sinh^2x=\left[\frac{1}{2}(e^x-e^{-x})\right]^2\]

zepdrix (zepdrix):

\[\Large\rm \frac{1}{4}(e^x-e^{-x})^2\]Understand how to expand that binomial out?

OpenStudy (anonymous):

no really. my teacher gave me an example and it has my head swimming.

zepdrix (zepdrix):

So recall how we expand a square binomial:\[\Large\rm (a-b)^2=a^2-2ab+b^2\]It's the square of the first, the square of the second, and also 2 times the product of both.

zepdrix (zepdrix):

\[\Large\rm (e^x-e^{-x})^2=(e^x)^2-2e^x e^{-x}+(e^{-x})^2\]

zepdrix (zepdrix):

Then we need to use rules of exponents to simplify the squares. Recall your exponent rule:\[\Large\rm (a^b)^c=a^{bc}\]

OpenStudy (anonymous):

okay